solve and check
One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers

Answer by ankor@dixie-net.com(22740)

"One number is 6 more than another" translates to:
one number = x; another number = (x+6)
:
If the sum of the smaller number and 3 times the larger number is 34,
x + 3(x+6) = 34
:
find the two numbers.
x + 3(x+3) = 34
x + 3x + 18 = 34
4x = 34 - 18
4x = 16
x = 16/4
x = 4
:
Obvious "another" number is; 4 + 6 = 10